Recent Log Entries

iCalibrate: Hall A

Log Number: 4130344Logbook(s): Body: Calibration ParametersHall:ATarget: no targetPass: 4Note: Calibrated On:2023-Feb-08 01:15Calibrated By: kpriceFile:A-2023-02-08_0115_no_target_new_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton1,875</

Fast SEE FTT at 8E

Log Number: 4130341Logbook(s): Entry Makers: carlinoBody: 

Not sure exactly how to interpret this, but seems a wide band of signals at the 50 - 100 Hz regime.

iCalibrate: Hall C

Log Number: 4130324Logbook(s): Body: Calibration ParametersHall:CTarget: Solid Thick Target (Gold)Pass: 5Note: Used for all foilsCalibrated On:2023-Jan-12 03:43Calibrated By: kiarraFile:C-2023-01-12_0343_Solid_Thick_Target__Gold__Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

Fopt

Log Number: 4130321Logbook(s): Tags: Body: Fopt 7.7 completed successfully. The runtime settings include a sample count of 15, a minimum current of 4, and an operating regime of Standard with Excitation in 1S Region. The corrector and RF kicks applied were 1.48165 and 1.48165 volts respectively. The max. amplitude setting was 1 millimeters. The files for this run were saved in the directory /cs/opshome/hlapps/optics/diforb/23feb131209. The X and Y stability of the machine during the run was 0.0551324 and 0.0685151 (mm) respectively.

HYST AREA 6-2-4 Cycles Magnets

Log Number: 4130320Logbook(s): Body: 

HYST AREA Version 6-2-4 has successfully cycled the following:

Area.Halls.HallA
Area.Halls.HallB
Area.Halls.HallC
Area.Halls.HallD.Extraction
Area.Halls.HallD.HallD

Total Time: 8.167 mins

GP-PL

Log Number: 4130319Logbook(s): Entry Makers: schoeneBody: 

GP adjusted, PL checked and not changed.

iocsl26rfint restored

Log Number: 4130316Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2023-02-13 11:22:33 The output from the myRestore execution was as follows: iocsl26rfint Rf12Intlk_Analog@2023-02-13 11:22 55(55) restored Rf12Intlk_Slope@2023-02-13 11:22 121(121) restored Pause: 5 second(s) Execute: caput R2QXIARCinit 1 (success) Old : R2QXIARCinit 0 New : R2QXIARCinit 1 system@2023-02-13 11:22 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsl26rfint (success) Procedure return status: Success

iocsl26rffcc4 restored

Log Number: 4130315Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2023-02-13 11:22:28 The output from the myRestore execution was as follows: iocsl26rffcc4 Rf12Fcc_V2 Pause: 60 second(s) Execute: /cs/opshome/burt/REBOOT/Rf12Fcc/DLOAD_ICAL.csh iocsl26rffcc4 (success) c100_rffcc@2023-02-13 11:22 32(32) restored Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-ecal/iocsl26rffcc4-ecal.snap (success) c100_ecal@2023-02-13 11:22 19(19) restored Execute: myRestore -execute -P c100_ecal -t R2Q4_ECAL -m BLDGNUM=2 ZONE=Q CAV=4 (success) c100_ecal@2022-09

HYST AREA 6-2-4 Cycles Magnets

Log Number: 4130312Logbook(s): Body: 

HYST AREA Version 6-2-4 has successfully cycled the following:

Area.Linacs.North
Area.Linacs.South
Area.Arcs.Arc1.Spreader
Area.Arcs.Arc1.Arc
Area.Arcs.Arc1.Recombiner
Area.Arcs.Arc1.Extraction
Area.Arcs.Arc2.Spreader
Area.Arcs.Arc2.Arc
Area.Arcs.Arc2.Recombiner
Area.Arcs.Arc2.Extraction
Area.Arcs.Arc3.Spreader
Area.Arcs.Arc3.Arc
Area.Arcs.Arc3.Recombiner
Area.Arcs.Arc3.Extraction
Area.Arcs.Arc4.Spreader
Area.Arcs.Arc4.Arc
Area.Arcs.Arc4.Recombiner
Area.Arcs.Arc4.Extraction

LEM South Linac

Log Number: 4130305Logbook(s): Tags: Body: LEM 9.15 was successfully run for the South Linac on Mon Feb 13 11:14:35 2023. The application phase took 0:00:14 to complete.

FudgeIt

Log Number: 4130304Logbook(s): Tags: Body: 

FudgeIt 4.4 changed the South Linac fudge factor from 1.001 to 0.997.

Pages